Please Answer A-BYour company has two divisions: One division sells software and the other division sells computers through a direct sales cha
Answer #1

Given,

Dell's beta = 1.21

Risk free rate = 4.7%

Cost of debt = 6.2%

Tax rate = 38% or 0.38

Market risk premium = 5.8%

Solution :-

Cost of equity for Dell = Rf tab (market disk Bemium) Where, Rf - disk-free date b = beta of Dell Now, Cost of equity for Delweight of Dells Debt (ud) = 1 - weight of Dells equity = 1 - 0.98979 = 0.00021 W ACC for Dell = (ke (We) + (Kd) (Wa) where,After tax Cost of debt (ka). = Cost of debt (l- tax rate) 6.2% (1-038) 2 6.2% (0.62) = 3.844% Now, W ACC for del = (17.718% (weight Sales of Computer division (Wo) = 40% g 0.40 weight of - 1-0.40 software division (Ws) = 0.60 Now, Overall company WAC8.144% 0.60 – WACC of software division 13.57% = WACC of software division Thus, Estimate of the WACC for your software divis